OpenStudy (jackfrost33):

>Divide the following polynomial, then place the answer in the proper location on the grid. Write your answer in order of descending powers of y. (y 3 - 4y 2 + 7y - 6) (y - 2)

OpenStudy (mww):

find out how many times y goes into y^3 (forget about the -2 part for now and the other powers) *Then write your answer on top as y^2 Then multiply the divisor (y-2) by this. y^2(y-2)= y^3 - 2y^2 Write this below then subtract this from your dividend (first two terms) (y^3 - 4y^2) - (y^3 -2y^2) = -2y^2 and write it below, and drag down the next power (7y). *Divide -2y^2 by y again to get -2y, then multiply (y-2) by this and subtract. -2y(y-2) = -2y^2 + 4y -2y^2+7y - (-2y^2 + 4y) = 3y. Drag down the next power (6) and *divide 3y by y to get 3, multiply, subtract... 3(y-2) = 3y - 6 which when subtracted from 3y - 6 gives 0. So we have fully divided now. Answer to the division will be y^2 - 2y +3